# Minimum numbers of Dehn colors of knots and $\mathcal{R}$-palette graphs

### v1

@@ 5. Evaluations of ${\rm mincol}^{\rm Dehn}_{p}(K)$ by $\mathcal{R}$-palette graphs
Let $S\subset\mathbb{Z}_{p}$. If $S=\mathcal{C}(D,C)$ for some nontrivially Dehn $p$-colored diagram $(D,C)$ of a knot, the $\mathcal{R}$-palette graph $G_{S}$ includes a connected $\mathcal{R}$-subgraph with at least three vertices.
@@
From the assumption, the $\mathcal{R}$-palette graph $G_{S}$ includes the $\mathcal{R}$-palette graph $G_{(D,C)}=(V_{(D,C)},E_{(D,C)})$ as an $\mathcal{R}$-subgraph. Hence we will show that $G_{(D,C)}$ is connected, and $G_{(D,C)}$ includes at least three vertices.
First, let us discuss the connectedness of $G_{(D,C)}$. For any $\mathbin{\overline{b}},\mathbin{\overline{b^{\prime}}}\in V_{(D,C)}$ with $\mathbin{\overline{b}}\not=\mathbin{\overline{b^{\prime}}}$, there exist two arcs $u$ and $u^{\prime}$ of $D$ such that $u$ and $u^{\prime}$ are a $\mathbin{\overline{b}}$-arc and a $\mathbin{\overline{b^{\prime}}}$-arc, respectively. Since $D$ is a knot diagram, there exists a finite sequence
$u=u_{1},c_{1},u_{2},c_{2},\ldots,c_{n-1},u_{n}=u^{\prime}$
of arcs $u_{1},\ldots,u_{n}$ and crossings $c_{1},\ldots,c_{n-1}$ of $D$ such that $u_{i}$ and $u_{i+1}$ are the under-arcs of $c_{i}$ for each $i\in\{1,\ldots n-1\}$. This implies that there exists a finite sequence
$\mathbin{\overline{b}}=\mathbin{\overline{b_{1}}},e_{1},\mathbin{\overline{b_{2}}},e_{2},\ldots,e_{n-1},\mathbin{\overline{b_{n}}}=\mathbin{\overline{b^{\prime}}}$
of $\mathbin{\overline{b_{1}}},\mathbin{\overline{b_{2}}},\ldots,\mathbin{\overline{b_{n}}}\in V_{(D,C)}$ and $e_{1},e_{2},\ldots,e_{n-1}\in E_{(D,C)}$ such that $u_{i}$ is a $\mathbin{\overline{b_{i}}}$-arc for each $i\in\{1,\ldots n\}$, and $e_{i}$ is the edge between $\mathbin{\overline{b_{i}}}$ and $\mathbin{\overline{b_{i+1}}}$ for each $i\in\{1,\ldots n-1\}$, where if $\mathbin{\overline{b_{i}}}=\mathbin{\overline{b_{i+1}}}$, then $e_{i}=\emptyset$. Thus, $\mathbin{\overline{b}}$ and $\mathbin{\overline{b^{\prime}}}$ can be connected.
Next, let us discuss the reason that $G_{(D,C)}$ includes at least three vertices. Since $C$ is nontrivial, $(D,C)$ includes at least one nontrivially colored crossing $c$ as depicted in the left, the center or the right of Figure 7, where $a_{1},\ldots,a_{4}$ are distinct elements of $\mathbb{Z}_{p}$. For the case of the crossing in the left of Figure 7, we have the $\mathbin{\overline{2a_{1}}}$-, $\mathbin{\overline{a_{1}+a_{2}}}$-, and $\mathbin{\overline{2a_{2}}}$-arcs whose labels are all distinct. For the case of the crossing in the center of Figure 7, we have the $\mathbin{\overline{a_{1}+a_{2}}}$-, $\mathbin{\overline{2a_{1}}}$-, and $\mathbin{\overline{a_{1}+a_{3}}}$-arcs whose labels are all distinct. For the case of the crossing in the right of Figure 7, we have the $\mathbin{\overline{a_{1}+a_{2}}}$-, $\mathbin{\overline{a_{1}+a_{3}}}$-, and $\mathbin{\overline{a_{3}+a_{4}}}$-arcs whose labels are all distinct. Thus we have at least three vertices of $G_{(D,C)}$. This completes the proof. ∎
Figure 7.
@@
For any subset $S\subset\mathbb{Z}_{7}$ such that #$\#S\leq 3$, the $\mathcal{R}$-palette graph $G_{S}$ includes no connected $\mathcal{R}$-subgraph with at least three vertices. This and Theorem 5.1 imply that ${\rm mincol}^{\rm Dehn}_{7}(K)\geq 4$ for any Dehn $7$-colorable knot $K$.
Any subset of $\mathbb{Z}_{7}$ with four elements is affine equivalent to either $\{0,1,2,3\}$ or $\{0,1,2,4\}$. We note that by Lemma 1.2 and Remark 1.3, when we select candidates of sets of four colors each of which might be the set of colors for a nontrivially Dehn $p$-colored diagram of a knot, it suffices to discuss only the representatives of the affine equivalence classes, that is $\{0,1,2,3\}$ and $\{0,1,2,4\}$. As shown in the right of Figure 4, the $\mathcal{R}$-palette graph $G_{\{0,1,2,4\}}$ clearly has a connected $\mathcal{R}$-subgraph with at least three vertices, which is $G_{\{0,1,2,4\}}$ itself for example. On the other hand, the $\mathcal{R}$-palette graph $G_{\{0,1,2,3\}}$ depicted in the left of Figure 4 does not include a connected $\mathcal{R}$-subgraph with at least three vertices. Therefore, if there exists a nontrivially Dehn $7$-colored diagram $(D,C)$ for a knot with #$\#\mathcal{C}(D,C)=4(=\lfloor\log_{2}7\rfloor+2)$, then $\mathcal{C}(D,C)$ is affine equivalent to $\{0,1,2,4\}$, and moreover, there exists a Dehn $7$-coloring $C^{\prime}$ of $D$ with $C\sim C^{\prime}$ and $\mathcal{C}(D,C^{\prime})=\{0,1,2,4\}$.
@@
    There is no subset $S\subset\mathbb{Z}_{p}$ with at most $\lfloor\log_{2}p\rfloor+1$ elements such that the $\mathcal{R}$-palette graph $G_{S}$ includes a connected $\mathcal{R}$-subgraph with at least three vertices. This implies that
@@
This can be shown by Mathematica computations. In particular, we found a connected $\mathcal{R}$-subgraph with at least three vertices for the $\mathcal{R}$-palette graph of each candidate of the sets of colors given in (2) of this theorem as depicted in Figures 812. ∎
@@
For $p=3$ (resp. $5,7,11,17,19,23,31$), the diagram illustrated in Figure 14 (resp. 14, 16, 16, 18, 18, 20, 20) admits the Dehn $p$-coloring with $\lfloor\log_{2}p\rfloor+2$ colors. Hence, from these properties and Theorem 1.5, we can see that there exists a knot $K$ with ${\rm mincol}^{\rm Dehn}_{p}(K)=\lfloor\log_{2}p\rfloor+2$ for each $p\in\{3,5,7,11,17,19,23,31\}$.
@@
from $b\not=c$, $a\not=b$; $a\not=2b+1$, and $b\not=4$, respectively, and the 2-gons with no label in the right of Figures 30 and 31 are not colored by $4$.
@@ Acknowledgments
The authors wish to express their thanks to Professor Shin Satoh for several helpful comments on the property ${\rm mincol}^{\rm Dehn}_{5}(K)=4$. The second author was supported by JSPS KAKENHI Grant Number 21K03233.

### v2

@@ 4. $\mathcal{R}$-palette graphs
We remark that $\bar{b}$ for a $\bar{b}$-arc was defined just a symbol, but $\bar{b}$ is regarded as an equivalence class as in Definition 4.1 when it represents a vertex.
We also remark that this $\mathcal{R}$-palette graph of $(D,C)$ is isomorphic to the palette graph of $(D,\bar{C})$ defined in [8], where $\bar{C}$ is the Fox $p$-coloring corresponding to $C$ as shown in Remark 1.1. Furthermore, it was shown in [8] that the palette graph of $(D,\bar{C})$ is connected and has at least $\lfloor\log_{2}p\rfloor+2$ vertices.
@@ 5. Evaluations of ${\rm mincol}^{\rm Dehn}_{p}(K)$ by $\mathcal{R}$-palette graphs
Let $S\subset\mathbb{Z}_{p}$. If $S=\mathcal{C}(D,C)$ for some nontrivially Dehn $p$-colored diagram $(D,C)$ of a knot, the $\mathcal{R}$-palette graph $G_{S}$ includes a connected $\mathcal{R}$-subgraph with at least $\lfloor\log_{2}p\rfloor+2$ vertices.
@@
This is evident because $G_{S}$ includes $G_{(D,C)}$ and as remarked in the previous section, $G_{(D,C)}$ is connected and has at least $\lfloor\log_{2}p\rfloor+2$ vertices. ∎
@@
For any subset $S\subset\mathbb{Z}_{7}$ such that #$\#S\leq 3$, the $\mathcal{R}$-palette graph $G_{S}$ includes no connected $\mathcal{R}$-subgraph with at least $\lfloor\log_{2}7\rfloor+2$ vertices. This and Theorem 5.1 imply that ${\rm mincol}^{\rm Dehn}_{7}(K)\geq 4$ for any Dehn $7$-colorable knot $K$.
Any subset of $\mathbb{Z}_{7}$ with four elements is affine equivalent to either $\{0,1,2,3\}$ or $\{0,1,2,4\}$. We note that by Lemma 1.2 and Remark 1.3, when we select candidates of sets of four colors each of which might be the set of colors for a nontrivially Dehn $p$-colored diagram of a knot, it suffices to discuss only the representatives of the affine equivalence classes, that is $\{0,1,2,3\}$ and $\{0,1,2,4\}$. As shown in the right of Figure 4, the $\mathcal{R}$-palette graph $G_{\{0,1,2,4\}}$ clearly has a connected $\mathcal{R}$-subgraph with at least $\lfloor\log_{2}7\rfloor+2$ vertices, which is $G_{\{0,1,2,4\}}$ itself for example. On the other hand, the $\mathcal{R}$-palette graph $G_{\{0,1,2,3\}}$ depicted in the left of Figure 4 does not include a connected $\mathcal{R}$-subgraph with at least $\lfloor\log_{2}7\rfloor+2$ vertices. Therefore, if there exists a nontrivially Dehn $7$-colored diagram $(D,C)$ for a knot with #$\#\mathcal{C}(D,C)=4(=\lfloor\log_{2}7\rfloor+2)$, then $\mathcal{C}(D,C)$ is affine equivalent to $\{0,1,2,4\}$, and moreover, there exists a Dehn $7$-coloring $C^{\prime}$ of $D$ with $C\sim C^{\prime}$ and $\mathcal{C}(D,C^{\prime})=\{0,1,2,4\}$.
@@
    There is no subset $S\subset\mathbb{Z}_{p}$ with at most $\lfloor\log_{2}p\rfloor+1$ elements such that the $\mathcal{R}$-palette graph $G_{S}$ includes a connected $\mathcal{R}$-subgraph with at least $\lfloor\log_{2}p\rfloor+2$ vertices. This implies that
@@
This can be shown by Mathematica computations. In particular, we found a connected $\mathcal{R}$-subgraph with at least $\lfloor\log_{2}p\rfloor+2$ vertices for the $\mathcal{R}$-palette graph of each candidate of the sets of colors given in (2) of this theorem as depicted in Figures 711. ∎
@@
For $p=3$ (resp. $5,7,11,17,19,23,31$), the diagram illustrated in Figure 13 (resp. 13, 15, 15, 17, 17, 19, 19) admits the Dehn $p$-coloring with $\lfloor\log_{2}p\rfloor+2$ colors. Hence, from these properties and Theorem 1.5, we can see that there exists a knot $K$ with ${\rm mincol}^{\rm Dehn}_{p}(K)=\lfloor\log_{2}p\rfloor+2$ for each $p\in\{3,5,7,11,17,19,23,31\}$.
@@
from $b\not=c$, $a\not=b$; $a\not=2b+1$, and $b\not=4$, respectively, and the 2-gons with no label in the right of Figures 29 and 30 are not colored by $4$.
@@ Acknowledgments
The authors wish to express their thanks to Professor Shin Satoh for several helpful comments on Theorem 5.1 and the property ${\rm mincol}^{\rm Dehn}_{5}(K)=4$. The second author was supported by JSPS KAKENHI Grant Number 21K03233.